Complete the table below.

A. Source charge (Q) B. Distance from source charge (R) C. Electric potential (Ve)

(A.1) 4.31x10^3 C (B.1) 7.11x10^-12m (C.1) ________________?

(A.2) ____________? (B.2) 3.05x10^-6m (C.2) 3.32V

(A.3) 2.33x10^-13C (B.3) _____________? (C.3) 12.04V

(A.4) 2.04x10^9C (B.4) 5.83x10^13m (C.4)________________?



1
Expert's answer
2021-03-21T11:19:57-0400

"(C.1)\ V_e=k\frac{q}{r}=9\cdot10^9\cdot\frac{4.31\cdot10^3}{7.11\cdot10^{-12}}=5.46\cdot10^{24}\ (V)"


"(A.2)\ q=\frac{V_e\cdot r}{k}=\frac{3.32\cdot3.05\cdot10^{-6}}{9\cdot10^9}=1.03\cdot10^{-15}\ (C)"


"(B.3)\ r=k\frac{q}{V_e}=9\cdot10^9\cdot\frac{2.33\cdot10^{-13}}{12.04}=1.74\cdot10^{-4}\ (m)"


"(C.4)\ V_e=k\frac{q}{r}=9\cdot10^9\cdot\frac{2.04\cdot10^9}{5.83\cdot10^{13}}\approx315\cdot10^{3}\ (V)"